# Measurement-Induced Spectral Transition

### v1

@@ Abstract
While we have focused on the spectral gap, we next compare the entire spectrum of the monitored dynamics with that of Hamiltonians in one-dimensional systems. Spectra of locally interacting many-body Hamiltonians usually have the following two structures: (a) the width of the spectrum becomes $O(L)$, which means that the typical level spacing is $O(e^{-L})$, and (b) the spectral gap becomes $O[1/\mathrm{poly}(L)]$ in gapless regimes or constant above a few states exponentially close to the ground state in gapped phases. We can analytically show that $\{\varepsilon_{i}(\eta)\}$ of our monitored dynamics also satisfy (a). Indeed, employing the majorization, the Horn theorem, and properties of doubly stochastic matrices [49], we can obtain [44]
@@
On the other hand, the monitored dynamics do not satisfy (b) in the gapless phase; the Lyapunov spectra of low-lying excited states become $O(e^{-L})$, which is numerically illustrated in Supplemental Material [44].
Discussion.–
We have shown that the quantum dynamics exposed to weak measurements exhibit the spectral transition in terms of the Lyapunov spectrum, whose transition point coincides with that of the entanglement transition. This coincidence is analogous to ground-state phase transitions between gapless and gapped phases, while the scalings of the entanglement and the gap in the monitored dynamics are qualitatively different from those in isolated quantum systems. To establish the analogy and clarify the distinction between these transitions, exploring the locality of $K_{t}(\eta)$ should be an important future direction. While ground states in condensed matter models are usually determined from local Hamiltonians, it is non-trivial whether $K_{t}(\eta)$ becomes local. This distinction may be the origin of the difference in the scalings of the spectral gap and entanglement in gapless phases of these two transitions.
We have also shown that the spectral transition corresponds to the transition of the timescale for the memory loss of a pure state and conjectured that it also corresponds to the purification transition of mixed states. In monitored dynamics with projective measurements, where the purification transition is also reported [22], singular values of time-evolution operators can be strictly zero, and the Lyapunov spectrum with $i>1$ become ill-defined when the projective measurement is done to all qubits. In contrast, singular values $\{\Lambda_{t,i}(\eta)\}$ do not become zero in weak measurements, which may make relaxation dynamics under weak measurements distinct from those under projective measurements. Exploring the discrepancy and similarity between projective and weak measurements should be an interesting future work.

### v2

@@ Abstract
While we have focused on the spectral gap, we next compare the entire spectra of our effective Hamiltonian and Hamiltonians in one-dimensional systems. Spectra of locally interacting many-body Hamiltonians usually have the following two structures: (a) the width of the spectrum scales at most linearly with $L$, which means that the typical level spacing is $\exp[-O(L)]$, and (b) the spectral gap becomes $O[1/\mathrm{poly}(L)]$ in gapless regimes or constant above a few states exponentially close to the ground state in gapped phases. We can analytically show that $\{\varepsilon_{i}(\eta)\}$ of our monitored dynamics also satisfy (a). Indeed, employing the majorization, the Horn theorem, and properties of doubly stochastic matrices [54], we can obtain [47]
@@
On the other hand, the monitored dynamics do not satisfy (b) in the gapless phase; the Lyapunov spectra of low-lying excited states become $\exp[-O(L)]$, which is numerically illustrated in Supplemental Material [47].
Conclusion.–
We have shown that the quantum dynamics exposed to generalized measurements exhibit the spectral transition, whose threshold coincides with that of the entanglement transition. We have also shown that the spectral transition leads to the transition of the memory-loss timescale of pure states and argued that it also leads to the purification transition of mixed states.
Our results suggest that Lyapunov analysis can reveal different measurement-induced phases in analogy with equilibrium quantum phase transitions. It is thus an intriguing future work to explore other counterparts of equilibrium phase transitions, such as topological transitions and spontaneous symmetry breaking, in monitored dynamics via Lyapunov analysis. It is also important to study the relation between our effective Hamiltonian and the Hamiltonians in replica methods [20, 33, 41].